OpenStudy (anonymous):

Assuming that a,b, and a/b are whole numbers and b does not equal 0 , explain which properties, definitions, or theorems would be used to show (a/b) x b = a. Then use them in your explanation.

OpenStudy (anonymous):

bcoz a/b is awhole number let a/b=k(say) then a= bk= kb =(a/b) *b hope that is a simple ans

OpenStudy (anonymous):

I am still a little bit confused, is their a particular property, definition, or theorem that you could possible direct me too, so I can maybe look it up? thanks

OpenStudy (anonymous):

bcoz a/b is a whole number then a/b must be an integer and thus it follows.